# Enhancing Biomedical Knowledge Discovery for Diseases: An Open-Source Framework Applied on Rett Syndrome and Alzheimer’s Disease

### v2

@@ Appendix I Probing: Additional Experiments
  • Entity Check: Examine the entities and their types. If an entity is incorrect, if the entity span is inaccurate (includes irrelevant words), or if the entity type is incorrect (e.g., "Rett syndrome" categorized as part of the human body), click the "Remove First Entity" or "Remove Second Entity" button, corresponding to the error.
  • Removing a Sentence: If a sentence lacks informative content, you have the option to remove it. Use this option if you are confident the sentence is uninformative.
@@
  • Negative Relation: The two entities are negatively correlated. This is a rare case, and negative words or phrases (e.g., "no," "absence") often indicate this.
@@
    Important: Once you press "Done", the instance can’t be retrieved, so ensure your decision is accurate.
@@
    Provide Relation Context: First, you need to finalize your choice for the relation labeling and then provide (if any) the related piece of text. If classifying a pair as related, specify the word or phrase in the sentence that influenced your decision. Use the text box provided and preferably copy-paste to avoid spelling errors. Press "Enter" after inputting the text to store it.

### v3

@@ Appendix H Detailed Annotation Guidelines
  • Entity Check: Examine the entities and their types. If an entity is incorrect, if the entity span is inaccurate (including irrelevant words), or if the entity type is incorrect (e.g., ”Rett syndrome” categorized as part of the human body), click the ”Remove First Entity” or ”Remove Second Entity” button, corresponding to the error.
  • Removing a Sentence: If a sentence lacks informative content, you have the option to remove it. Use this option if you are confident that the sentence is uninformative.
@@
  • Negative Relation: The two entities are negatively correlated. This is a rare case, and negative words or phrases (e.g., ”no,” ”absence”) often indicate this.
@@
    Important: Once you press ”Done”, the instance can’t be retrieved, so ensure your decision is accurate.
@@
    Provide Relation Context: First, you need to finalize your choice for the relation labeling and then provide (if any) the related piece of text. If classifying a pair as related, specify the word or phrase in the sentence that influenced your decision. Use the text box provided and preferably copy-paste to avoid spelling errors. Press ”Enter” after inputting the text to store it.